### v2

# CoDefeater: Using LLMs To Find Defeaters in Assurance Cases

Abstract.
Constructing assurance cases is a widely used, and sometimes required, process toward demonstrating that safety-critical systems will operate safely in their planned environment. To mitigate the risk of errors and missing edge cases, the concept of defeaters - arguments or evidence that challenge claims in an assurance case - has been introduced. Defeaters can provide timely detection of weaknesses in the arguments, prompting further investigation and timely mitigations. However, capturing defeaters relies on expert judgment, experience, and creativity and must be done iteratively due to evolving requirements and regulations. This new ideas paper proposes CoDefeater, an automated process to leverage large language models (LLMs) for finding defeaters. Initial results on two systems show that LLMs can efficiently find known and unforeseen feasible defeaters to support safety analysts in enhancing the completeness and confidence of assurance cases.
1. Introduction
Safety-critical systems have become deeply integrated into many societal domains, including healthcare, transportation, energy, and aviation (48; 22; 33). Failures in these systems can lead to catastrophic consequences for human safety, including fatalities and environmental and property damage (61). This has led to an increased focus on their dependence, reliability, and safety (59; 40). Many systems must comply with regulations (56; 10), provide evidence of safety, and undergo rigorous certification processes (36; 55) for approval from regulatory bodies. Assurance cases (ACs) have emerged as a common practice for this purpose, facilitating the verification of system correctness and the validation of specific claims regarding safety, security, and trustworthiness, among others (47; 8; 6).
An assurance case is a structured hierarchy of claims and arguments supported by evidence that a system will function as intended in a specified environment (62; 8). Several formal notations (e.g., Goal Structuring Notation (GSN) (37), Claims-Arguments-Evidence (CAE) (2), and Eliminative Argumentation (EA) (29)), along with tools (21; 65; 47), have been proposed. However, concerns arise over their completeness, uncertainty, and soundness for cyber-physical systems (31; 25), leading to false confidence and catastrophic failures (64). For example, failure of the minimum safe altitude warning system that led to a major aviation accident was attributed to incomplete and flawed reasoning in the safety case (32).
To enhance the robustness of assurance cases, it is critical to identify and mitigate their defeaters (also known as assurance weakeners). Defeaters highlight gaps in evidence or reasoning that undermine the validity of claims in the assurance case (34). An example of a defeater, drawn from the assurance case for the safe operation of an sUAS (small Uncrewed Aircraft System) battery, challenges the assurance case’s claim that “The sUAS has enough battery charge to complete its mission.” The defeater casting doubt on this claim is, “Unless the battery monitor is not calibrated/inaccurate.” The defeater serves to record, within the assurance case itself, the analyst’s challenge to the validity of the claim.
Various approaches have been proposed to identify and mitigate defeaters (30; 38). However, manually creating defeaters is a labor-intensive and time-consuming process (53; 50), relying heavily on safety analysts’ judgment, experience, creativity, and understanding of the system. This can lead to confirmation bias (65; 7). As assurance cases evolve with new standards and technological advances, ongoing efforts are focused on formal and semi-automated approaches for detecting and managing defeaters (23; 42). Best practices build assurance cases incrementally, so automating all/or part of this process is important (68; 23).
Large Language Models (LLMs) are increasingly automating software engineering tasks like test generation and defect detection (35; 3). In particular, these models have become valuable in tasks requiring complex understanding, such as vulnerability detection, requirements elicitation, and code generation (69; 17; 5). Moreover, LLMs have shown promise in automating evaluation tasks and acting as surrogate evaluators (18; 70). Consequently, we explore whether LLMs’ capabilities can be harnessed to automate defeater analysis toward the completeness, soundness, and confidence of assurance cases. Despite calls for more research into LLMs’ ability to identify defeaters (39; 66), no study has been conducted to evaluate and investigate their effectiveness.
This new ideas paper presents the first empirical investigation of the feasibility and utility of LLMs for identifying defeaters, using a process we call CoDefeater. It evaluates their potential to aid practitioners and safety analysts in the iterative human-in-the-loop process of finding defeaters, as shown in Figure 1. We evaluate the performance of an LLM (ChatGPT) in automated defeater analysis on two complex real-world case studies. Our experimental results suggest that CoDefeater is a promising approach for identifying and generating novel assurance case defeaters. Overall, this work makes three key contributions. 1) To the best of our knowledge, we provide the first empirical results from an investigation of the effectiveness and usefulness of an LLM (GPT 3.5) in identifying and creating defeaters for real-world assurance cases. 2) We provide a new assurance case fragment with defeaters that can be leveraged for further research on automated defeater identification techniques. 3). Based on our findings, we outline current challenges and directions for future work. All experimental artifacts are available here: https://gitlab.com/anonymousdot/codefeater.
2. Background and Related Work
Assurance case arguments typically adopt an inductive approach, where sub-claims offer direct evidence to support the parent claim but do not ensure it with certainty (8; 13). Focusing solely on proving a claim may introduce confirmation bias, as exemplified by the Nimrod aircraft crash (12). Recent approaches have therefore embraced defeasible reasoning, which acknowledges that arguments about system properties in practice are inherently defeasible (28; 44; 58). Defeaters are potential doubts or objections that challenge the validity of a claim, reflecting gaps in evidence and reasoning (28; 45; 25; 52). Figure 2 shows a fragment of an assurance case for an sUAS battery, with examples of defeaters (red boxes).
Defeaters are typically represented using EA notation (29) but have also been integrated into GSN and CAE notations (34; 8). Hence, this study does not aim to evaluate LLMs’ performance in identifying defeaters in any specific notation or semantic accuracy but rather to investigate the feasibility of this approach. Finally, the indefeasibility criterion requires a thorough search for defeaters in an assurance case (8), motivating our investigation into LLMs’ potential to assist practitioners and safety analysts in identifying and generating novel defeaters.
In software engineering, LLMs are increasingly employed to assist developers and automate tasks such as discovering requirements, code generation, testing, and program synthesis (35; 5). Diemert and Weber (24) have reported ChatGPT’s effectiveness in hazard analysis for safety-critical systems, highlighting their potential to assist human analysts. In the context of software assurance cases, Sivakumar et al. (63) assessed LLM’s (GPT-4) proficiency in understanding GSN representations and its performance in constructing safety cases. Viger et al. (66) proposed using LLMs to identify defeaters; however, they did not report an empirical evaluation of their capabilities. Khakzad Shahandashti et al. (39) explored LLM’s (GPT-4) understanding of EA notation and defeater concepts, with empirical validation left as future work. We aim to address this gap by examining LLM performance in identifying defeaters for assurance cases in two case studies.
The Machine Learning (ML) community also has explored the reasoning abilities of LLMs (71; 20; 67). This line of inquiry investigates how LLMs fare in open-ended tasks and their effectiveness in assisting humans. Findings have indicated that LLMs can demonstrate consistent responses exhibiting similarities with human evaluations, suggesting their potential as automated tools (18).
3. Methodology
We conducted a preliminary exploratory study toward answering the following two research questions:
RQ1: (Effectiveness). How effective are LLMs in identifying and analyzing defeaters in assurance cases?
RQ2: (Utility). Can LLMs support practitioners in generating novel and meaningful defeaters?
3.1. Experimental Setup
Datasets. We performed our experiments on two assurance cases. The first is for the CERN Large Hadron Collider (LHC) Machine Protection System (MPS), which provides assurance that the MPS will prevent damage to the LHC from unstable, high-energy particle beams (60; 51). The assurance case for the CERN LHC uses EA notation, and we extracted claim nodes with their corresponding defeaters for our experiments. The second assurance case is a fragment of a larger one that our team has recently created for small Uncrewed Aircraft Systems (sUAS)111Anonymized for submission using GSN. It addresses the claim “the sUAS has enough charge in its battery to complete the mission” (see claim 1.2 in Figure 2). We include both the assurance case fragment and corresponding defeaters in the supplementary material222https://gitlab.com/anonymousdot/codefeater. The real-world complexity of these systems and the availability of defeaters made them suitable for our preliminary experiments. Table 1 provides a summary of the assurance cases.
Model. We used ChatGPT (GPT-3.5) for our study, specifically model GPT-3.5-turbo released by Open AI (11).
| Node Type | LHC MPS | sUAS Battery |
| Claims | 61 | 6 |
| Defeaters | 103 | 20 |
3.2. Prompt Design
An effective prompt design is crucial for achieving good performance, as the choice of prompts significantly impacts the quality, relevance, and accuracy of the LLM’s response (15). It involves crafting a system prompt to establish the context and prepare the LLM for the task, along with a user prompt that contains the specific task request (54). We designed the system prompt following OpenAI’s best practices (54) and relevant literature in software engineering (72; 14) and open-ended evaluation tasks (18). Our process identified role-based system prompts (43) as the most effective approach. Figure 3 shows the system prompt used in our study.
Several user prompting techniques have been proposed, e.g., zero-shot, one-shot, few-shot, and chain-of-thought (15). Zero-shot learning involves providing the model with only the task description (system prompt), without examples of unseen tasks to learn from. In contrast, one-shot and few-shot learning conditions the model on one or more examples in the prompt, respectively (11).
For our preliminary study, we adopted the zero-shot learning setting. This approach both (1) facilitates the immediate, off-the-shelf application of LLMs, eliminating the need for computationally expensive fine-tuning procedures, and (2) is naturally suited to scenarios such as ours, with limited data availability for training or fine-tuning (41). Each prompt was presented independently to the model to avoid influencing subsequent responses, allowing us to assess its standalone capabilities (16).
3.3. Evaluation Criteria
Due to the complexity and open-ended nature of the task (e.g., varying response length vs. ground truth, and subjectivity), automatic evaluation metrics were not suitable (18). Therefore, we relied on human evaluation for assessing LLM performance. For RQ1, we used a deductive coding approach (46; 24), where the first two authors independently reviewed responses and categorized them as complete match, partial match, or no match based on similarity to ground-truth defeaters. We used the defeaters in the LHC assurance case as the ground truth. For the battery assurance case, a set of ground-truth defeaters was provided by one of the authors (independently) familiar with the domain, following best practices (51). For RQ2, the responses were evaluated for being reasonable (14), i.e., the defeater could reasonably be in the ground truth but had been overlooked. This aimed to assess the LLM’s capability to identify novel defeaters. Figure 4 shows an example of each type of match. Next, the reviewers met to discuss and finalize their assigned codes. In the case of post-discussion disagreement, if one reviewer labeled a response as a partial match and the other as a complete match, we categorized it as a partial match to avoid confirmation bias (26). In the one instance where one reviewer indicated no match while the other identified a partial or complete match, it was discarded. Last, we calculated inter-rater agreement using Cohen’s Kappa (19) to evaluate the consistency and reliability of the coding process.
3.4. Threats to Validity
There is potential subjectivity in the qualitative evaluation of LLM performance on defeater identification. To address that, two authors independently coded the LLM responses, following best practices (9; 27), and held multiple discussions to avoid misinterpretations. We also computed Cohen’s kappa (19), which indicated substantial inter-reviewer agreement. To avoid confirmation bias, disagreements were coded as partial or no match. The non-deterministic nature of LLMs and different versions might produce slightly different responses; however, we used a single version of ChatGPT. Finally, the preliminary results presented here lead us to propose that the use of LLMs to generate defeaters merits further work; however, generalizability awaits larger studies with improved LLMs.
4. Results
In this section, we present the key findings of our experiments, grouped by our research questions.
4.1. (RQ1): Effectiveness in Identifying Defeaters
[style = exampledefault] Finding 1: The LLM displayed promising zero-shot capabilities for defeater analysis in assurance cases.
Figure 5 presents the results of our closed coding process, which assessed the GPT-3.5 model’s performance in identifying defeaters for the two real-world assurance cases in a zero-shot setting. Our experiments showed that the model demonstrated unexpectedly good zero-shot capabilities for defeater analysis. Specifically, it completely identified more than half of all defeaters and partially identified more than a third in both datasets. Fewer than of the defeaters created by human analysts were not identified at all by the LLM. These results were noteworthy given the complexity of the systems under study and the lack of system information, domain knowledge, or few-shot examples provided to the LLM. Additionally, the total coded responses achieved high inter-rater agreement, with a Cohen’s Kappa score of 0.81, (19; 49). Finally, all of these defeaters were identified in the first prompt, illustrating the convenience of the approach. Among the unidentified defeaters, some required specific domain or system knowledge not provided to the model, indicating areas for potential improvement.
[style = exampledefault] Finding 2: The LLM struggled with defeaters that challenged implicit assumptions.
We conducted a manual analysis of the unidentified defeaters (n = 17) to investigate whether there were any patterns behind the LLM’s failure. Interestingly, we found that the model struggled to identify those defeaters that implicitly challenged the truth of an assumption. For example, for the claim ”The BICs will not produce a FALSE BEAM-PERMIT to trigger a beam dump, unless a loss of the high-frequency signal (10Mhz) in either Beam Permit Loop (A and B) is detected...”, a (ground-truth) defeater in the LHC assurance case questions the assumption that the 10Mhz signal is the right signal to monitor. In other words, the presence of another similar high-frequency signal might lead to a false indication of TRUE BEAM-PERMIT. Unlike the analysts, the LLM did not question the underlying assumption in the claim and thus did not identify the defeater. Consequently, future work should explore integrating external knowledge sources since, in similar tasks, it significantly enhances LLM’s performance (72; 57).
4.2. (RQ2): Utility in Generating Novel Defeaters
To answer RQ2, we evaluated the LLM’s performance on the sUAS battery’s assurance case, where we have the necessary domain knowledge. Using the same prompting method, we iteratively requested additional defeaters to assess its capacity to generate novel defeaters beyond the ones that had been identified in the building of the assurance case.
[style = exampledefault] Finding 3: LLMs can support practitioners in providing useful and novel defeaters.
The LLM output was a useful source of five novel defeaters, each of which was feasible upon further investigation using our evaluation criteria. These were: (1) an unexpected power drain due to an onboard component failure; (2) an emergency external to the sUAS that forced the sUAS into a longer flight; (3) a missed waypoint to which the pilot had to return; (4) unexpected power drain arising from ongoing efforts to recover a lost GPS; and (5) external interference that the sUAS had to dodge repeatedly. The last one was interesting to us because the LLM gave as an example of interference that birds might attack the sUAS. This, in fact, happens quite often and is dangerous (4). The response shows how an LLM can offer a creative perspective that catches missing edge cases.
5. Discussion
Based on our findings, we highlight several challenges and opportunities for an LLM-based process to help find defeaters.
Designing better prompts. Prompt designing has been shown to significantly impact LLM performance (43; 15). Many prompting methods have been proposed, and further investigation for suitability to defeater analysis is needed. Moreover, our study revealed that the LLM can generate creative, redundant, and far-fetched scenarios (e.g., defeaters due to budget constraints). Balancing LLM creativity with defeater relevance poses an important challenge.
Rationale behind defeaters. In our experiments, we found that the LLM responses not only identified defeaters but also provided helpful rationale and examples. For instance, if the ground truth defeater stated, ”Unless there are incorrect readings,” the LLM suggested, ”The sensors may not be properly calibrated, leading to inaccurate readings.” These explanations can assist analysts in understanding and analyzing both a defeater’s feasibility and its potential mitigations. Investigating explainable prompting techniques such as Chain-Of-Thought (15) is an important next step.
Towards incremental assurance using LLMs. Our study focused on single claims and associated defeaters. Future research should evaluate the performance of LLMs on a combination of claims. It will be interesting to investigate whether LLMs can identify the impact of defeaters on multiple claims and assess if the provided evidence adequately addresses them. This direction will require developing detailed data for evidence analysis and exploring prompts specifically designed for this purpose.
6. Conclusion and Future Work
We have presented CoDefeater, a process for automated defeater discovery in assurance cases using LLMs (GPT-3.5). Our evaluation on two real-world case studies demonstrated the LLM’s zero-shot capabilities in identifying defeaters and its potential to support practitioners in an iterative human-in-the-loop process. We make available the portion of a new assurance case and its ground-truth defeaters used in our experiments for other researchers. Future work will expand beyond the zero-shot setting to explore one-shot and few-shot learning approaches for improved performance. Additionally, fine-tuning LLMs on assurance cases presents an avenue to improve their performance. Our study provides preliminary results as a starting point for future research to explore the role of LLMs as a tool to assist with the identification of defeaters toward the development of improved assurance cases.
Acknowledgements.
This work was funded by grant 80NSSC23M0058 from the National Aeronautics and Space Administration (NASA).References
- (1)
- Adelard (2024) Adelard. 2024. https://www.adelard.com/asce/cae/
- Ahmed and Devanbu (2022) Toufique Ahmed and Premkumar Devanbu. 2022. Few-shot training llms for project-specific code-summarization. In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering. 1–5.
- Alliance (2024) Golden Gate Bird Alliance. 2024. https://goldengatebirdalliance.org/conservation/make-the-city-safe-for-wildlife/drone-dangers-and-birds/
- Arora et al. (2024) Chetan Arora, John Grundy, and Mohamed Abdelrazek. 2024. Advancing requirements engineering through generative ai: Assessing the role of llms. In Generative AI for Effective Software Development. Springer, 129–148.
- Assurance Case Working Group (2021) Assurance Case Working Group. 2021. Assurance Case Guidance. Technical Report SCSC-159.
- Bloomfield et al. (2021) Robin Bloomfield, Gareth Fletcher, Heidy Khlaaf, Luke Hinde, and Philippa Ryan. 2021. Safety case templates for autonomous systems. arXiv preprint arXiv:2102.02625 (2021).
- Bloomfield and Rushby (2020) Robin Bloomfield and John Rushby. 2020. Assurance 2.0: A manifesto. arXiv preprint arXiv:2004.10474 (2020).
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Breaux et al. (2006) Travis D Breaux, Matthew W Vail, and Annie I Anton. 2006. Towards regulatory compliance: Extracting rights and obligations to align requirements with regulations. In 14th IEEE International Requirements Engineering Conference (RE’06). IEEE, 49–58.
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems 33 (2020), 1877–1901.
- Cave (2006) CH Cave. 2006. An independent review into the broader issues surrounding the loss of the RAF Nimrod MR2 Aircraft XV230 in Afghanistan in 2006. The Stationary Office, Tech. Rep (2006).
- Chechik (2019) Marsha Chechik. 2019. Uncertain requirements, assurance and machine learning. In 2019 IEEE 27th International Requirements Engineering Conference (RE). IEEE, 2–3.
- Chen et al. (2023a) Boqi Chen, Kua Chen, Shabnam Hassani, Yujing Yang, Daniel Amyot, Lysanne Lessard, Gunter Mussbacher, Mehrdad Sabetzadeh, and Dániel Varró. 2023a. On the use of GPT-4 for creating goal models: an exploratory study. In 2023 IEEE 31st International Requirements Engineering Conference Workshops (REW). IEEE, 262–271.
- Chen et al. (2023c) Banghao Chen, Zhaofeng Zhang, Nicolas Langrené, and Shengxin Zhu. 2023c. Unleashing the potential of prompt engineering in large language models: a comprehensive review. arXiv preprint arXiv:2310.14735 (2023).
- Chen et al. (2023b) Lingjiao Chen, Matei Zaharia, and James Y. Zou. 2023b. How is ChatGPT’s behavior changing over time? ArXiv abs/2307.09009 (2023). https://api.semanticscholar.org/CorpusID:259951081
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. 2021. Evaluating large language models trained on code. arXiv preprint arXiv:2107.03374 (2021).
- Chiang and Lee (2023) Cheng-Han Chiang and Hung-yi Lee. 2023. Can large language models be an alternative to human evaluations? arXiv preprint arXiv:2305.01937 (2023).
- Cohen (1960) Jacob Cohen. 1960. A coefficient of agreement for nominal scales. Educational and psychological measurement 20, 1 (1960), 37–46.
- Dai et al. (2023) Shih-Chieh Dai, Aiping Xiong, and Lun-Wei Ku. 2023. LLM-in-the-loop: Leveraging large language model for thematic analysis. arXiv preprint arXiv:2310.15100 (2023).
- Denney et al. (2012) Ewen Denney, Ganesh Pai, and Josef Pohl. 2012. AdvoCATE: An assurance case automation toolset. In Computer Safety, Reliability, and Security: SAFECOMP 2012 Workshops: Sassur, ASCoMS, DESEC4LCCI, ERCIM/EWICS, IWDE, Magdeburg, Germany, September 25-28, 2012. Proceedings 31. Springer, 8–21.
- Dey et al. (2018) Nilanjan Dey, Amira S Ashour, Fuqian Shi, Simon James Fong, and João Manuel RS Tavares. 2018. Medical cyber-physical systems: A survey. Journal of medical systems 42 (2018), 1–13.
- Diemert et al. (2023) Simon Diemert, John Goodenough, Jeff Joyce, and Charles Weinstock. 2023. Incremental assurance through eliminative argumentation. Journal of System Safety 58, 1 (2023), 7–15.
- Diemert and Weber (2023) Simon Diemert and Jens H Weber. 2023. Can Large Language Models Assist in Hazard Analysis?. In International Conference on Computer Safety, Reliability, and Security. Springer, 410–422.
- Duan et al. (2017) Lian Duan, Sanjai Rayadurgam, Mats PE Heimdahl, Anaheed Ayoub, Oleg Sokolsky, and Insup Lee. 2017. Reasoning about confidence and uncertainty in assurance cases: A survey. In Software Engineering in Health Care: 4th International Symposium, FHIES 2014, and 6th International Workshop, SEHC 2014, Washington, DC, USA, July 17-18, 2014, Revised Selected Papers 4. Springer, 64–80.
- Gemalmaz and Yin (2021) Meric Altug Gemalmaz and Ming Yin. 2021. Accounting for Confirmation Bias in Crowdsourced Label Aggregation.. In IJCAI. 1729–1735.
- Gohar et al. (2024) Usman Gohar, Michael C Hunter, Agnieszka Marczak-Czajka, Robyn R Lutz, Myra B Cohen, and Jane Cleland-Huang. 2024. Towards Engineering Fair and Equitable Software Systems for Managing Low-Altitude Airspace Authorizations. arXiv preprint arXiv:2401.07353 (2024).
- Goodenough et al. (2012) J. Goodenough, Charles B. Weinstock, and Ari Z. Klein. 2012. Toward a Theory of Assurance Case Confidence. https://api.semanticscholar.org/CorpusID:13431021
- Goodenough et al. (2015) John B Goodenough, Charles B Weinstock, and Ari Z Klein. 2015. Eliminative argumentation: A basis for arguing confidence in system properties. Software Engineering Institute, Carnegie Mellon University, Pittsburgh, PA, Tech. Rep. CMU/SEI-2015-TR-005 (2015).
- Graydon and Holloway (2017) Patrick J Graydon and C Michael Holloway. 2017. An investigation of proposed techniques for quantifying confidence in assurance arguments. Safety science 92 (2017), 53–65.
- Greenwell et al. (2004a) William S. Greenwell, C. Michael Holloway, and John C. Knight. 2004a. A Taxonomy of Fallacies in System Safety Arguments. https://api.semanticscholar.org/CorpusID:267911368
- Greenwell et al. (2004b) William S Greenwell, Elisabeth A Strunk, and John C Knight. 2004b. Failure analysis and the safety-case lifecycle. In Human Error, Safety and Systems Development: IFIP 18th World Computer Congress TC13/WC13. 5 7th Working Conference on Human Error, Safety and Systems Development 22–27 August 2004 Toulouse, France. Springer, 163–176.
- Gunes et al. (2014) Volkan Gunes, Steffen Peter, Tony Givargis, and Frank Vahid. 2014. A survey on concepts, applications, and challenges in cyber-physical systems. KSII Transactions on Internet and Information Systems (TIIS) 8, 12 (2014), 4242–4268.
- Hawkins et al. (2011) Richard Hawkins, Tim Kelly, John Knight, and Patrick Graydon. 2011. A new approach to creating clear safety arguments. In Advances in Systems Safety: Proceedings of the Nineteenth Safety-Critical Systems Symposium, Southampton, UK, 8-10th February 2011. Springer, 3–23.
- Hou et al. (2023) Xinying Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John C. Grundy, and Haoyu Wang. 2023. Large Language Models for Software Engineering: A Systematic Literature Review. ArXiv abs/2308.10620 (2023). https://api.semanticscholar.org/CorpusID:261048648
- Johnson et al. (1998) Leslie A Johnson et al. 1998. DO-178B: Software considerations in airborne systems and equipment certification. Crosstalk, October 199 (1998), 11–20.
- Kelly and Weaver (2004) Tim Kelly and Rob Weaver. 2004. The goal structuring notation–a safety argument notation. In Proceedings of the dependable systems and networks 2004 workshop on assurance cases, Vol. 6. Citeseer Princeton, NJ.
- Khakzad Shahandashti et al. (2023) Kimya Khakzad Shahandashti, Alvine Boaye Belle, Timothy Lethbridge, Oluwafemi Odu, and Mithila Sivakumar. 2023. A PRISMA-driven Systematic Review on System Assurance Weakeners.
- Khakzad Shahandashti et al. (2024) Kimya Khakzad Shahandashti, Mithila Sivakumar, Mohammad Mahdi Mohajer, Alvine Boaye Belle, Song Wang, and Timothy Lethbridge. 2024. Assessing the Impact of GPT-4 Turbo in Generating Defeaters for Assurance Cases. In Proceedings of the 2024 IEEE/ACM First International Conference on AI Foundation Models and Software Engineering (Lisbon, Portugal) (FORGE ’24). Association for Computing Machinery, New York, NY, USA, 52–56. https://doi.org/10.1145/3650105.3652291
- Knight (2002) John C Knight. 2002. Safety critical systems: challenges and directions. In Proceedings of the 24th international conference on software engineering. 547–550.
- Kojima et al. (2022) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2022. Large language models are zero-shot reasoners. Advances in neural information processing systems 35 (2022), 22199–22213.
- Kokaly et al. (2016) Sahar Kokaly, Rick Salay, Valentin Cassano, Tom Maibaum, and Marsha Chechik. 2016. A model management approach for assurance case reuse due to system evolution. In Proceedings of the ACM/IEEE 19th International Conference on Model Driven Engineering Languages and Systems (Saint-malo, France) (MODELS ’16). Association for Computing Machinery, New York, NY, USA, 196–206. https://doi.org/10.1145/2976767.2976792
- Kong et al. (2023) Aobo Kong, Shiwan Zhao, Hao Chen, Qicheng Li, Yong Qin, Ruiqi Sun, and Xin Zhou. 2023. Better zero-shot reasoning with role-play prompting. arXiv preprint arXiv:2308.07702 (2023).
- Koons (2008) Robert C. Koons. 2008. Defeasible Reasoning. In Stanford Encyclopedia of Philosophy.
- Koopman (2022) Philip Koopman. 2022. How Safe Is Safe Enough: Measuring and Predicting Autonomous Vehicle Safety. Carnegie Mellon University, https://books.google.com/books?id=lNVozwEACAAJ.
- Kulatunga et al. (2007) Udayangani Kulatunga, Dilanthi Amaratunga, and Richard Haigh. 2007. Structuring the unstructured data: the use of content analysis. (2007).
- Maksimov et al. (2018) Mike Maksimov, Nick LS Fung, Sahar Kokaly, and Marsha Chechik. 2018. Two decades of assurance case tools: a survey. In Computer Safety, Reliability, and Security: SAFECOMP 2018 Workshops, ASSURE, DECSoS, SASSUR, STRIVE, and WAISE, Västerås, Sweden, September 18, 2018, Proceedings 37. Springer, 49–59.
- Maksimov et al. (2019) Mike Maksimov, Sahar Kokaly, and Marsha Chechik. 2019. A Survey of Tool-supported Assurance Case Assessment Techniques. ACM Comput. Surv. 52, 5, Article 101 (sep 2019), 34 pages. https://doi.org/10.1145/3342481
- McHugh (2012) Mary L McHugh. 2012. Interrater reliability: the kappa statistic. Biochemia medica 22, 3 (2012), 276–282.
- Menghi et al. (2023) Claudio Menghi, Torin Viger, Alessio Di Sandro, Chris Rees, Jeff Joyce, and Marsha Chechik. 2023. Assurance Case Development as Data: A Manifesto. In 45th IEEE/ACM International Conference on Software Engineering: New Ideas and Emerging Results, NIER@ICSE, Melbourne, Australia, May 14-20, 2023. IEEE, 135–139.
- Millet et al. (2023) Laure Millet, Simon Diemert, Chris Rees, Torin Viger, Marsha Chechik, Claudio Menghi, and Jeffrey Joyce. 2023. Assurance Case Arguments in the Large: The CERN LHC Machine Protection System. In International Conference on Computer Safety, Reliability, and Security. Springer, 3–10.
- Murugesan et al. (2023) Anitha Murugesan, Isaac Hong Wong, Robert J. Stroud, Joaquín Arias, Elmer Salazar, Gopal Gupta, Robin Bloomfield, Srivatsan Varadarajan, and John Rushby. 2023. Semantic Analysis of Assurance Cases using s(CASP). In Proceedings of the International Conference on Logic Programming 2023 Workshops co-located with the 39th International Conference on Logic Programming (ICLP 2023), London, United Kingdom, July 9th and 10th, 2023 (CEUR Workshop Proceedings, Vol. 3437). CEUR-WS.org.
- Nair et al. (2014) Sunil Nair, Jose Luis De La Vara, Mehrdad Sabetzadeh, and Lionel Briand. 2014. An extended systematic literature review on provision of evidence for safety certification. Information and Software Technology 56, 7 (2014), 689–717.
- OpenAI (2024) OpenAI. 2024. Retrieved 2024 from https://platform.openai.com/docs/guides/prompt-engineering
- Palin et al. (2011) Rob Palin, David Ward, Ibrahim Habli, and Roger Rivett. 2011. ISO 26262 safety cases: Compliance and assurance. (2011).
- Parliament and Council (2016) European Parliament and E Council. 2016. General data protection regulation. official Journal of the European Union 59 (2016), 294.
- Peng et al. (2023) Baolin Peng, Michel Galley, Pengcheng He, Hao Cheng, Yujia Xie, Yu Hu, Qiuyuan Huang, Lars Liden, Zhou Yu, Weizhu Chen, et al. 2023. Check your facts and try again: Improving large language models with external knowledge and automated feedback. arXiv preprint arXiv:2302.12813 (2023).
- Pollock (1987) John L Pollock. 1987. Defeasible reasoning. Cognitive science 11, 4 (1987), 481–518.
- Rausand (2014) Marvin Rausand. 2014. Reliability of safety-critical systems: theory and applications. John Wiley & Sons.
- Rees et al. (2023) Chris Rees, Rolf Lippelt, Marsha Chechik, Lukas Felsberger, Mateo Delgado, Markus Zerlauth, Jeff Joyce, Claudio Menghi, Jan Uythoven, Simon Diemert, and et al. 2023. Assessing the usefulness of assurance cases: An experience with the CERN Large Hadron Collider. http://cds.cern.ch/record/2854725
- Rushby (1994) John Rushby. 1994. Critical system properties: Survey and taxonomy. Reliability Engineering & System Safety 43, 2 (1994), 189–219.
- Rushby (2015) John Rushby. 2015. The interpretation and evaluation of assurance cases. Comp. Science Laboratory, SRI International, Tech. Rep. SRI-CSL-15-01 (2015).
- Sivakumar et al. (2023) Mithila Sivakumar, Alvine Boaye Belle, Jinjun Shan, and Kimya Khakzad Shahandashti. 2023. GPT-4 and Safety Case Generation: An Exploratory Analysis. arXiv preprint arXiv:2312.05696 (2023).
- UL Muram et al. (2018) Faiz UL Muram, Barbara Gallina, and Laura Gómez Rodríguez. 2018. Preventing Omission of Key Evidence Fallacy in Process-Based Argumentations. In 2018 11th International Conference on the Quality of Information and Communications Technology (QUATIC). 65–73. https://doi.org/10.1109/QUATIC.2018.00019
- Varadarajan et al. (2023) Srivatsan Varadarajan, Robin Bloomfield, John Rushby, Gopal Gupta, Anitha Murugesan, Robert Stroud, Kateryna Netkachova, and Isaac Hong Wong. 2023. CLARISSA: Foundations, Tools & Automation for Assurance Cases. In 2023 IEEE/AIAA 42nd Digital Avionics Systems Conference (DASC). 1–10. https://doi.org/10.1109/DASC58513.2023.10311213
- Viger et al. (2023) Torin Viger, Logan Murphy, Simon Diemert, Claudio Menghi, Alessio Di, and Marsha Chechik. 2023. Supporting Assurance Case Development Using Generative AI. In SAFECOMP 2023, Position Paper. Toulouse, France. https://hal.science/hal-04191791
- Wang et al. (2024) Yiqi Wang, Wentao Chen, Xiaotian Han, Xudong Lin, Haiteng Zhao, Yongfei Liu, Bohan Zhai, Jianbo Yuan, Quanzeng You, and Hongxia Yang. 2024. Exploring the Reasoning Abilities of Multimodal Large Language Models (MLLMs): A Comprehensive Survey on Emerging Trends in Multimodal Reasoning. ArXiv abs/2401.06805 (2024). https://api.semanticscholar.org/CorpusID:266999728
- Warg et al. (2019) Fredrik Warg, Hans Blom, Jonas Borg, and Rolf Johansson. 2019. Continuous deployment for dependable systems with continuous assurance cases. In 2019 IEEE International Symposium on Software Reliability Engineering Workshops (ISSREW). IEEE, 318–325.
- Zhang et al. (2023) Chenyuan Zhang, Hao Liu, Jiutian Zeng, Kejing Yang, Yuhong Li, and Hui Li. 2023. Prompt-enhanced software vulnerability detection using chatgpt. arXiv preprint arXiv:2308.12697 (2023).
- Zheng et al. (2024) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al. 2024. Judging llm-as-a-judge with mt-bench and chatbot arena. Advances in Neural Information Processing Systems 36 (2024).
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric P. Xing, Haotong Zhang, Joseph Gonzalez, and Ion Stoica. 2023. Judging LLM-as-a-judge with MT-Bench and Chatbot Arena. ArXiv abs/2306.05685 (2023). https://api.semanticscholar.org/CorpusID:259129398
- Zhou et al. (2024) Xin Zhou, Ting Zhang, and David Lo. 2024. Large language model for vulnerability detection: Emerging results and future directions. In Proceedings of the 2024 ACM/IEEE 44th International Conference on Software Engineering: New Ideas and Emerging Results. 47–51.